# Hierarchical Procedural Framework for Low-latency Robot-Assisted Hand-Object Interaction

### v2

@@ Abstract
Advances in robotics have been driving the development of human-robot interaction (HRI) technologies. However, accurately perceiving human actions and achieving adaptive control remains a challenge in facilitating seamless coordination between human and robotic movements. In this paper, we propose a hierarchical procedural framework to enable dynamic robot-assisted hand-object interaction. An open-loop hierarchy leverages the computer vision (CV)-based 3D reconstruction of the human hand, based on which motion primitives have been designed to translate hand motions into robotic actions. The low-level coordination hierarchy fine-tunes the robot’s action by using the continuously updated 3D hand models. Experimental validation demonstrates the effectiveness of the hierarchical control architecture. The adaptive coordination between human and robot behavior has achieved a delay of $\leq 0.3$ seconds in the tele-interaction scenario. A case study of ring-wearing tasks indicates the potential application of this work in assistive technologies such as healthcare and manufacturing.
@@ 1 INTRODUCTION
To physically assist the human partners, these robots/agents are equipped with sensing and manipulation capabilities such that the co-bots can detect, interpret, and infer human cues, thereby providing tangible support to the human. To date, one typical task to achieve sophisticated pHRI has spotlighted the hand-object interaction (HOI), where humans manipulate objects and interact with the environments [7, 8]. Recent developments in computer vision have shed light on the robot-assisted HOI, such as using depth-enhanced vision systems and soft-robotic gloves to predict and assist with various hand postures for rehabilitation [9].
There are two fundamental technical challenges for robot-assisted HOI. One is hand pose estimation, which aims to accurately detect, track, and reconstruct the human hand in a dynamic environment under perturbation [10]. Research has focused on 3D hand reconstruction using diverse technologies such as wearable sensors, radio frequency (RF) technologies, and acoustic methods. In particular, the development of 3D pose estimation based on single RGB images has become prevalent in the current HOI community due to its ease of deployment. For instance, [11] leveraged the graph convolutional neural network to reconstruct the whole 3D mesh of the hand surface from a single RGB image, which can well represent hand shape variations and capture local details. [12] introduced a contact potential field approach, treating interactions as a spring-mass system to allow for dynamic adjustments based on contact physics, which enhances the physical plausibility of HOI predictions, even in cases of significant discrepancies in ground-truth data. In contrast, [13] integrated the transformer architecture with graph convolution to create the MeshGraphormer, which adeptly captures both local and global interactions among mesh vertices and body joints. In summary, the computer vision-based approaches provide a cost-effective and end-to-end solution for capturing intricate hand movements that advance the design of pHRI platforms.
@@
Another challenge lies in designing the robotic controller for robot-assisted manipulation. Various efforts have been put toward developing control interfaces, such as virtual joysticks [14], electromyography [15], inertial measurement units (IMUs) [16], eye gaze detection [17] and brain-machine interfaces [18]. Notably, motion primitives (MPs)-based control has emerged as a promising method, which decomposes complex tasks into smaller, reusable actions that can be dynamically adapted. For instance, [19] introduced a hierarchical skill-learning framework that combines dynamic MPs with Gaussian mixture models to learn motor skills from human demonstrations. This approach allows robots to encode complex, human-like motions and adapt to various interaction scenarios by reproducing these skills during cooperative tasks. Similarly, [20] developed an MPs-based navigation planner that leverages deep collision prediction to ensure safe navigation in dynamic environments without relying on explicit maps. This method integrates MPs with deep learning techniques to predict collision probabilities, enabling real-time decision-making for navigation tasks.
In this paper, we propose a hierarchical procedural system for pHRI in the context of robot-assisted dressing. The first hierarchy consists of hand recognition and tele-manipulation. We proposed an MPs-based robot controller to allow users to realize various customized motions. First, we introduced an RGB-based 3D hand pose estimation model to perform the real-time vibration-tolerant state estimation. We trained a model to associate the hand state with the predefined MPs library, which contains MPs of different levels. Regarding the low-level hierarchy, we proposed a closed-loop control architecture to ensure the co-bot adjusts its motion to assist human partners adaptively. Finally, we introduce the action alignment model that allows for the adaptive orientation of the end-effector. Experimental tests have validated the effectiveness of the proposed hierarchical control framework in HOI tasks. Our proposed human-robot collaborative model can pave the way for future development of robot-assisted healthcare and medical treatment services.
2 IMAGE-BASED 3D HAND POSE SENSING
@@
As shown in Figure 1(b), MeshGraphormer first leverages a pre-trained CNN to extract grid features and a global feature vector from the given RGB images. These features are then tokenized and fed to a multi-layer transformer-based encoder to output coarse mesh tokens, and a multi-layer perceptron (MLP) is utilized to perform upsampling to obtain full mesh tokens and mesh vertices. Finally, the 3D coordinates of hand joints are computed using a MANO model from [21]. Denote by $F_{\bm{\theta}}$ the model of MeshGraphormer with network parameters $\bm{\theta}$, we have
$(\bm{x},\bm{y},\bm{z})=F_{\bm{\theta}}\left(P(\bm{I})\right),$ (1)
where $\bm{I}$ is a raw RGB image, $\bm{x},\bm{y},\bm{z}$ are 3D coordinates vectors, and $P(\cdot)$ denotes the image preprocessing operation including resizing, cropping, normalizing, etc.
@@
In practice, data oscillation may occur due to slight hand shaking during real-time detection, leading to unstable localization. To address the problem, we leverage the following moving average filter to smooth the detection data. Additionally, we implement asynchronous processing pipelines to optimize our system responsiveness and minimize latency in real-time applications for image preprocessing and pose estimation. Moreover, the hand pose estimation model can be replaced by a body pose model in our system, allowing the system to be adapted for full-body interaction scenarios. This flexibility is essential in complex environments where understanding hand and body gestures can significantly enhance the robot’s ability to interpret human intentions accurately.
3 MOTION PRIMITIVES MODEL
@@
Equipped with the derived 3D hand information, we further propose a motion primitives model (MPM) that converts human hand motions into robotic actions in real time. By utilizing the 3D coordinates information of hand joints, our model interprets and translates specific hand movements into corresponding mechanical responses. In particular, we define a set of MPs of different levels, where low-level primitives handle basic actions such as grasping or releasing, which are directly mapped to robot actions without additional inputs. High-level primitives, on the other hand, process more complex gestures that may involve multiple sequential or parallel actions, integrating contextual data from the environment or other sensory feedback. This hierarchical structure allows the MPM to adaptively handle a wide range of human intentions, ensuring fluid, context-aware collaboration between the human and robot in dynamic tasks.
@@
We build the MPM using a bidirectional long short-term memory (LSTM) network [22], a type of recurrent neural network particularly suited for sequence prediction problems. This choice is critical as it enables the model to retain context over time, which is essential for interpreting continuous and dynamic hand gestures. The LSTM network is trained on a self-collected dataset of hand motion sequences to encompass a spectrum of potential gestures and their corresponding robotic actions. Through this training, the network learns the individual gestures and their transitions, ensuring a fluid and intuitive control flow. Denote by $M_{\bm{\phi}}$ the MPM with network parameters $\bm{\phi}$, we have
@@ 4.1 Experimental Setup
We utilize the UR5 manipulator as the robotic arm and a two-finger RobotiQ gripper as the end-effector to execute pick-and-place tasks. Hand pose recognition is facilitated by an RGB camera (Intel RealSense). The experimental setup is illustrated in Figure 2a.
@@
To control the UR5, we employed the real-time data exchange (RTDE) control interface to enable real-time manipulation of the tool center point (TCP) of the robot arm. Motor servoing was achieved using the servoL function at a frame rate of 500 Hz. As a safety precaution during implementation, we configured a safety zone around the end-effector to mitigate the risk of potential human injury.
@@ 4.2 Controller Architecture
The interaction between the hand and the object requires the robot to exhibit adaptive behaviors based on human input, necessitating hierarchical control over the robot arm. The gripper control operates as an open-loop system, enabling the gripper to execute pinching and releasing actions on the object. The UR5 TCP control relies on predefined motion primitives. Upon program initialization, the camera can capture hand movements and identify these primitives. When a motion primitive is detected, the robot executes predetermined incremental motions, allowing the user to adjust the spatial location of the TCP manually.
Closed-loop control occurs during the cooperation stage (Figure 2b). When the cooperation condition is triggered, the robot continuously approaches the human hand at a constant speed while tracking hand pose and receiving hand state information, such as finger angular deviation and fingertip location. Utilizing this data, the robot dynamically adjusts its vertical position and orientation to ensure successful object delivery. Throughout this process, the robot adapts both its spatial position and end-effector orientation to collaborate effectively with the human operator. The transition between these control modes is triggered by the detection of specific motion primitives, i.e., the ”Ring” primitive, signaling the robot to enter a cooperation mode. The entire control architecture is depicted in Figure 2.
@@ 4.3 Action Alignment
A crucial aspect is the coordination between the object and the human. When the gripper holds the object, the gripper-object system is considered a rigid body. When the gripper moves towards the human finger, the controller dynamically adjusts to accommodate finger bending and environmental perturbations. We propose an action alignment model to enable real-time adjustment of the robot’s end-effector pose.
Figure 3: The spatial coordinates alignment between the hand and object.
In Figure 3, we define the spatial coordinate of the object as $O_{o}=[X_{o},Y_{o},Z_{o}]$. The center is established at the object’s center, with the x-axis aligned to the object’s normal direction. The y-axis corresponds to the gripper’s central line, while the z-axis follows the right-hand rule. Similarly, with the employed hand pose recognition model, we determine the fingertip’s coordinate as $O_{h}=[X_{h},Y_{h},Z_{h}]$. The hand coordinate system originates from the fingertip, with the y-axis aligned with the fingertip’s direction. The x-axis is positioned perpendicular to the y-axis and towards the bending side. The z-axis is set based on the right-hand rule. For the two coordinate systems defined in this initial system, it is always possible to find a rotation matrix T that performs the transformation from $O_{o}$ to the $O_{h}$:
@@
To ensure that the ring can be worn on the finger, the orientation should follows:
@@
The subscript indicates the coordinate system being referenced, while the superscript indicates the object (ring or finger). The symbols $v_{h}$ and $v_{v}$ represent the robot’s velocity in the horizontal and vertical directions, respectively. To ensure the alignment, we need to have $\alpha=-1$ and $\beta=1$ such that the robot’s TCP synchronizes with the human hand’s vertical movements, and when finger bending occurs, the TCP adjusts to align the object with the fingertip’s central axis.
@@ 5.1 Hand Recognition Performance
Figure 4: 3D modeling of the process of converting from the motion ”Keep” to the motion ”Come”. (a) Mesh reconstruction of the hand motion, in which the filled circles represent the joints. (b) Changes in the coordinates of joints during the motion conversion process. Here, ”IJ1-4” denotes the four joints of the index finger, respectively.
@@
We first evaluate the effectiveness of the designed 3D hand pose module. Table 1 illustrates the pre-defined motion primitives while Figure 4 demonstrates the 3D modeling of the transition process between two distinct hand motions. In the ”Keep” motion, the y-axis coordinates of the four joints remain relatively unchanged while under the ”Come” motion the coordinates of these joints progress according to the hand motion. These observations validate the capture and model of the hand motion changes, providing reliable reference information for the subsequent robot control.
@@ 5.2.1 Data Preparation
For each motion, the human participants were instructed to make and maintain the motion and then collected 2000 consecutive frames of RGB image data, and the subject made slight motion adjustments during the collection process. Then, these frames were used to extract 3D coordinates of 21 hand joints using the aforementioned pose estimation model. For a single trajectory, we sampled every 10 frames and intercepted a sequence of length 10. As a result, a motion dataset with 1000 samples was obtained, and the date shape of each input is $(10,63)$. Furthermore, we use stratified sampling to divide the data set into a training set (80%) and a validation set (20%) to ensure that each type of action has the same number of samples.
@@ 5.3.1 Motion Primitives Transition
Each motion primitive is activated individually. Supporting Video S1 demonstrates the motion control using the MPM, where the UR5 robot’s movements along the Y-axis are primarily controlled by two MPs. The robot exhibits a strong dynamic response to these commands, with different MPs triggering actions such as forward and backward movements. Figure 5a illustrates the time delays between the robot’s actions and the human commands during two consecutive come-back MPM commands. It can be seen that at each timestamp, the robot consistently lags behind the human command. For the ’come’ action, the delay is 0.168 seconds, while the ’back’ MPM action experiences a delay of approximately 0.24 seconds. The maximum control delay remains under 0.3 seconds, demonstrating the cobot’s effective command-tracking capabilities within the control architecture.

### v3

@@ Abstract
Advances in robotics have been driving the development of human-robot interaction (HRI) technologies. However, accurately perceiving human actions and achieving adaptive control remains a challenge in facilitating seamless coordination between human and robotic movements. In this paper, we propose a hierarchical procedural framework to enable dynamic robot-assisted hand-object interaction (HOI). An open-loop hierarchy leverages the RGB-based 3D reconstruction of the human hand, based on which motion primitives have been designed to translate hand motions into robotic actions. The low-level coordination hierarchy fine-tunes the robot’s action by using the continuously updated 3D hand models. Experimental validation demonstrates the effectiveness of the hierarchical control architecture. The adaptive coordination between human and robot behavior has achieved a delay of $\leq 0.3$ seconds in the tele-interaction scenario. A case study of ring-wearing tasks indicates the potential application of this work in assistive technologies such as healthcare and manufacturing.
@@ 1 INTRODUCTION
To physically assist the human partners, these robots/agents are equipped with sensing and manipulation capabilities such that the co-bots can detect, interpret, and infer human cues, thereby providing tangible support to the human. To date, one typical task to achieve sophisticated pHRI has spotlighted the hand-object interaction (HOI), where humans manipulate objects and interact with the environment [7, 8]. Recent developments in computer vision have shed light on the robot-assisted HOI, such as using depth-enhanced vision systems and soft-robotic gloves to predict and assist with various hand postures for rehabilitation [9].
There are two fundamental technical challenges for robot-assisted HOI. One is hand pose estimation, which aims to accurately detect, track, and reconstruct the human hand in a dynamic environment under perturbation [10]. Research has focused on 3D hand reconstruction using diverse technologies such as wearable sensors, radio frequency (RF) technologies, and acoustic methods [11]. In particular, the development of 3D pose estimation based on single RGB images has become prevalent in the current HOI community due to its ease of deployment. For instance, [12] leveraged the graph convolutional neural network to reconstruct the whole 3D mesh of the hand surface from a single RGB image, which can well represent hand shape variations and capture local details. [13] introduced a contact potential field approach, treating interactions as a spring-mass system to allow for dynamic adjustments based on contact physics, which enhances the physical plausibility of HOI predictions, even in cases of significant discrepancies in ground-truth data. In contrast, [14] integrated the transformer architecture with graph convolution to create the MeshGraphormer, which adeptly captures both local and global interactions among mesh vertices and body joints. In summary, computer vision-based approaches provide a cost-effective and end-to-end solution for capturing intricate hand movements that advance the design of pHRI platforms.
@@
Another challenge lies in designing the robotic controller for robot-assisted manipulation. Various efforts have been put toward developing control interfaces, such as virtual joysticks [15], electromyography [16], inertial measurement units (IMUs) [17], eye gaze detection [18] and brain-machine interfaces [19]. Notably, motion primitive (MP)-based control has emerged as a promising method, which decomposes complex tasks into smaller, reusable actions that can be dynamically adapted. For instance, [20] introduced a hierarchical skill-learning framework that combines dynamic MPs with Gaussian mixture models to learn motor skills from human demonstrations. This approach allows robots to encode complex, human-like motions and adapt to various interaction scenarios by reproducing these skills during cooperative tasks. Similarly, [21] developed an MP-based navigation planner that leverages deep collision prediction to ensure safe navigation in dynamic environments without relying on explicit maps. This method integrates MPs with deep learning techniques to predict collision probabilities, enabling real-time decision-making for navigation tasks.
In this paper, we propose a hierarchical procedural system for pHRI in the context of robot-assisted dressing. The first hierarchy consists of hand recognition and tele-manipulation. We proposed an MP-based robot controller to allow users to realize various customized motions. First, we introduce an RGB-based 3D hand pose estimation model to perform real-time vibration-tolerant state estimation. We trained a model to associate the hand state with MPs of different levels from a predefined MP library. Regarding the low-level hierarchy, we proposed a closed-loop control architecture to ensure the co-bot adjusts its motion to assist human partners adaptively. Finally, we introduce the action alignment model that allows for the adaptive orientation of the end-effector. Experimental tests have validated the effectiveness of the proposed hierarchical control framework in HOI tasks. Our proposed human-robot collaborative model contributes to the future development of robot-assisted healthcare and medical treatment services.
2 RGB-BASED 3D HAND POSE SENSING
@@
As shown in Figure 1(b), MeshGraphormer first leverages a pre-trained CNN to extract grid features and a global feature vector from the given RGB images. These features are then tokenized and fed to a multi-layer transformer-based encoder to output coarse mesh tokens, and a multi-layer perceptron (MLP) is utilized to perform upsampling to obtain full mesh tokens and mesh vertices. Finally, the 3D coordinates of hand joints are computed using a MANO model from [22]. Denote by $F_{\bm{\theta}}$ the model of MeshGraphormer with network parameters $\bm{\theta}$, we have $(\bm{x},\bm{y},\bm{z})=F_{\bm{\theta}}\left(P(\bm{I})\right)$, where $\bm{I}$ is a raw RGB image, $\bm{x},\bm{y},\bm{z}$ are 3D coordinates vectors, and $P(\cdot)$ denotes the image preprocessing operation including resizing, cropping, normalizing, etc.
@@
In practice, data oscillation may occur due to slight hand shaking during real-time detection, leading to unstable localization. To address the problem, we leverage a moving average filter to smooth the detection data. Additionally, we implement asynchronous processing pipelines to optimize our system’s responsiveness and minimize latency in real-time applications for image preprocessing and pose estimation. Moreover, the hand pose estimation model can be replaced by a body pose model in our system, allowing the system to be adapted for full-body interaction scenarios. This flexibility is essential in complex environments where understanding hand and body gestures can significantly enhance the robot’s ability to interpret human intentions accurately.
3 MOTION PRIMITIVE MODEL
@@
Equipped with the derived 3D hand information, we further propose a motion primitive model (MPM) that converts human hand motions into robotic actions in real time. By utilizing the 3D coordinates information of hand joints, our model interprets and translates specific hand motions into corresponding mechanical responses. In particular, we define a set of MPs of different levels, where low-level primitives handle basic actions such as grasping or releasing, which are directly mapped to robot actions without additional inputs. High-level primitives, on the other hand, process more complex gestures that may involve multiple sequential or parallel actions, integrating contextual data from the environment or other sensory feedback. This hierarchical structure allows the MPM to adaptively handle a wide range of human intentions, ensuring fluid, context-aware collaboration between the human and robot in dynamic tasks.
@@
We build the MPM using a bidirectional long short-term memory (LSTM) network [23], a type of recurrent neural network particularly suited for sequence prediction problems. This choice is important as it enables the model to retain context over time, which is essential for interpreting continuous and dynamic hand gestures. The LSTM network is trained on a self-collected dataset of hand motion sequences to encompass a spectrum of potential gestures and their corresponding robotic actions. Through this training, the network learns the individual gestures and their transitions, ensuring a fluid and intuitive control flow. Denote by $M_{\bm{\phi}}$ the MPM with network parameters $\bm{\phi}$, we have
@@ 4.1 Experimental Setup
We utilize the UR5 manipulator as the robotic arm and a two-finger RobotiQ gripper as the end-effector to execute pick-and-place tasks. Hand pose recognition is facilitated by an RGB camera (Intel RealSense). The experimental setup is illustrated in Figure 2(a).
@@
To control the UR5, we employed the real-time data exchange (RTDE) control interface to enable real-time manipulation of the tool center point (TCP) of the robot arm. Motor servoing was achieved using the servoL function at a control rate of 500 Hz. As a safety precaution during implementation, we configured a safety zone around the end-effector to mitigate the risk of potential human injury.
@@ 4.2 Controller Architecture
The interaction between the hand and the object requires the robot to exhibit adaptive behaviors based on human input, necessitating hierarchical control over the robot arm. The gripper control operates as an open-loop system, enabling the gripper to execute pinching and releasing actions on the object. The UR5 TCP control relies on predefined motion primitives. Upon program initialization, the camera can capture hand motion and identify these primitives. When a motion primitive is detected, the robot executes predetermined incremental motions, allowing the user to adjust the spatial location of the TCP manually.
Closed-loop control occurs during the cooperation stage (see Figure 2(b)). When the cooperation condition is triggered, the robot continuously approaches the human hand at a constant speed while tracking hand pose and receiving hand state information, such as finger angular deviation and fingertip location. Utilizing this data, the robot dynamically adjusts its vertical position and orientation to ensure successful object delivery. Throughout this process, the robot adapts both its spatial position and end-effector orientation to collaborate effectively with the human operator. The transition between these control modes is triggered by the detection of specific motion primitives, i.e., the ”Ring” primitive, signaling the robot to enter a cooperation mode. The entire control architecture is depicted in Figure 2.
@@ 4.3 Action Alignment
A crucial aspect is the coordination between the object and the human. When the gripper holds the object, the gripper-object system is considered a rigid body. When the gripper moves towards the human finger, the controller dynamically adjusts to accommodate finger bending and environmental perturbations. We design a motion alignment model to enable real-time adjustment of the robot’s end-effector pose.
Figure 3: The spatial coordinates alignment between the hand and the object.
In Figure 3, we define the spatial coordinate of the object as $O_{o}=[X_{o},Y_{o},Z_{o}]$. The center is established at the object’s center, with the x-axis aligned along the object’s normal direction. The y-axis corresponds to the gripper’s central line, while the z-axis follows the right-hand rule. Similarly, with the employed hand pose recognition model, we determine the fingertip’s coordinate as $O_{h}=[X_{h},Y_{h},Z_{h}]$. The hand coordinate system originates from the fingertip, with the y-axis aligned with the fingertip’s direction. The x-axis is positioned perpendicular to the y-axis and towards the bending side. The z-axis is set based on the right-hand rule. For the two coordinate systems defined in this initial system, it is always possible to find a rotation matrix T that performs the transformation from $O_{o}$ to the $O_{h}$:
@@
To ensure that the ring can be worn on the finger, the orientation should follow:
@@
The subscript indicates the coordinate system being referenced, while the superscript indicates the object (ring or finger). The symbols $v_{h}$ and $v_{v}$ represent the robot’s velocity in the horizontal and vertical directions, respectively. To ensure the alignment, we need to have $\alpha=-1$ and $\beta=1$ such that the robot’s TCP synchronizes with the human hand’s vertical movements, and when finger bending occurs, the TCP adjusts to align the object with the finger’s central axis.
@@ 5.1 Hand Recognition Performance
Figure 4: 3D modeling of the process of converting from the motion ”Keep” to the motion ”Come”. (a) Mesh reconstruction of the hand motion, in which the filled circles represent the joints. (b) Changes in the coordinates of joints during the motion conversion process. Here, ”IJ1-4” denotes the four joints of the index finger.
@@
We first evaluate the effectiveness of the designed 3D hand pose module. Table 1 illustrates the pre-defined motion primitives while Figure 4 demonstrates the 3D modeling of the transition process between two distinct hand motions. In the ”Keep” motion, the y-axis coordinates of the four joints remain relatively unchanged, while under the ”Come” motion, the coordinates of these joints progress according to the hand motion. These observations validate the capture and model of the hand motion changes, providing reliable reference information for the subsequent robot control.
@@ 5.2.1 Data Preparation
For each motion, the human participants were instructed to make and maintain the motion, and then collect 2000 consecutive frames of RGB image data, and the subject made slight motion adjustments during the collection process. Then, these frames were used to extract 3D coordinates of 21 hand joints using the aforementioned pose estimation model. For a single trajectory, we sampled every 10 frames and intercepted a sequence of length 10. As a result, a motion dataset with 1000 samples was obtained, and the data shape of each input is $(10,63)$. Furthermore, we use stratified sampling to divide the dataset into a training set (80%) and a validation set (20%) to ensure that each type of action has the same number of samples.
@@ 5.3.1 Motion Primitives Transition
Each motion primitive is activated individually. Supporting Video S1 demonstrates the motion control using the MPM, where the UR5 robot’s movements along the Y-axis are primarily controlled by two MPs. The robot exhibits a strong dynamic response to these commands, with different MPs triggering actions such as forward and backward movements. Figure 5(a) illustrates the time delays between the robot’s actions and the human commands during two consecutive come-back MPM commands. It can be seen that at each timestamp, the robot consistently lags behind the human command. For the ’come’ action, the delay is 0.168 seconds, while the ’back’ MPM action experiences a delay of approximately 0.24 seconds. The maximum control delay remains under 0.3 seconds, demonstrating the cobot’s effective command-tracking capabilities within the control architecture.